OpenStudy (anonymous):

What is the approximate distance from (4, 6) to (-2,1)? 5 units 9 units 6 units 8 units

OpenStudy (anonymous):

@HelpBlahBlahBlah @Hannnnah @annas @Loser66 @shrutipande9 @kirbykirby

OpenStudy (anonymous):

Please help!

OpenStudy (anonymous):

@OrangeMaster

OpenStudy (anonymous):

I can help

OpenStudy (anonymous):

use the distance formula \[d = \sqrt{(x2 - x1)^{2}+(y2 - y1)^2}\]

OpenStudy (anonymous):

appr. 8 units

OpenStudy (anonymous):

we have given (x1 , y1 ) =(4 , 6) & (x2 , y2) = (-2 , 1) So use above formula we get d = SQRT((-2 - 4)^ + (1-6)^2) = SQRT((-6)^2+(-5)^2) so d = SQRT(36 + 25) = SQRT(61) = appr 8 units